OpenStudy (anonymous):

Find the 7th partial sum of summation of 6 open parentheses 3 close parentheses to the I minus 1 power from 1 to infinity

OpenStudy (anonymous):

can you use math notation instead of words? \[\sum_1^\infty\]

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm \sum_{i=1}^{\infty}6(3)^{i-1}\]Is this what the problem should look like?

OpenStudy (anonymous):

yes it is

zepdrix (zepdrix):

I think the uhhh geometric series usually starts from an index of 0. So let's make the substitution \(\Large\rm n=i-1\). So where i starts at 1, n will start at 0 ( 1 - 1). \[\Large\rm \sum_{n=0}^{\infty}6(3)^{n}\]And so we have the form of a geometric series:\[\Large\rm \sum_{n=0}^{\infty}a r^n\]This series converges when \(\Large\rm r\lt1\). Does that hold true for our problem?

zepdrix (zepdrix):

Oh they wanted us to find the .... 7th partial sum..?

zepdrix (zepdrix):

Hmm I don't know what that is actually :(

zepdrix (zepdrix):

Does that just mean get the first 7 terms of the sum?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

I still think we should make this substitution (unless it's too confusing for you),\[\Large\rm \sum_{n=0}^{\infty}6(3)^{n}\]It just makes it easier to count. Starting from zero these would be our first 7 terms,\[\Large\rm =6(3)^0+6(3)^1+6(3)^2+6(3)^3+6(3)^4+6(3)^5+6(3)^6\]

zepdrix (zepdrix):

And from there, just use a calculator :U Understand how I plugged the values in for n?

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

and yes i did

OpenStudy (phi):

There is a formula you can use see https://en.wikipedia.org/wiki/Geometric_series#Formula btw, I think you want to sum from n=0 to n=6 to get the 7th partial sum the answer should be \[ 6 \frac{1-3^7}{1-3} \]